# Kummer-Artin-Schreier-Witt Theory

### v2

@@ 2.3.1 Deformed Artin-Hasse exponential function
The Artin-Hasse exponential series at $p$ is the formal power series
@@
$E_{s,p}(t):=\frac{E_{p}(\zeta_{p^{s}}t)}{E_{p}(t)}=\exp\left(\sum_{i=0}^{s-1}(\zeta_{p^{s-i}}-1)\frac{t^{p^{i}}}{p^{i}}\right)\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket t\rrbracket$ (11)
and, for $\underline{a}=(a_{0},\ldots,a_{s-1})$,
$E_{s,p}(\underline{a})=\prod_{i=0}^{s-1}E_{s-i,p}(a_{i})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket a_{0},\ldots,a_{s-1}\rrbracket.$ (12)
One shows that
@@
We define the formal power series
$G_{s,p}(\underline{a})=\frac{E_{s,p}(p\underline{a})}{E_{s-1,p}(\underline{a})}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket a_{0},\dots,a_{s-1}\rrbracket.$
We have
@@
We have $G_{s,p}(\underline{a})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket(\zeta_{p^{s}}-1)a_{0},\ldots,\lambda a_{s-1}\rrbracket$.
The series can be expressed in the below forms:
$G_{s,p}(\underline{a})$ $=\prod_{j=0}^{s-1}\exp\left(\sum_{i=0}^{s-j-1}(p(\zeta_{p^{s-j-i}}-1)-(\zeta_{p^{s-j-1-i}}-1))\frac{a_{j}^{p^{i}}}{p^{i}}\right)$
$=\prod_{j=0}^{s-1}\exp\left(\sum_{i=0}^{s-j-1}(\zeta_{p^{s-j-i}}-1)(p-1-\zeta_{p^{s-j-i}}-\cdots-\zeta_{p^{s-j-i}}^{p-1})\frac{a_{j}^{p^{i}}}{p^{i}}\right)$
$=\prod_{j=0}^{s-1}E_{s-j,p}\left((p-1-[\zeta_{p^{s-j}}]-\cdots-[\zeta_{p^{s-j}}^{p-1}])[a_{j}]\right),$
where $[a]=(a,0,\ldots,0)\in W_{s}(\mathbb{Z}_{(p)}[\zeta_{p^{s}}])$ is the Teichmüller lift of $a$.
@@
The $E_{i}$ (resp. $G_{i}$) from Theorem 1.1 are approximations of $E_{i,p}$ (resp. $G_{i,p}$).
2.3.2 Matsuda’s Kummer-Artin-Schreier-Witt theory
Let $K$ be a discrete valuation field containing $\zeta_{p^{s}}$, with residue field $k$. The Robba ring $\mathcal{R}$ over $K$ is defined as the ring of Laurent series that converge in some annulus $\{z\mid\rho<\lvert z\rvert<1\}$ for some $10<\rho<$ 1. Let $\mathcal{R}^{\rm bd}$ be the subring of $\mathcal{R}$ consisting of formal sums with bounded coefficients. The ring $\mathcal{R}^{\rm bd}$ carries a Gauss norm:
@@ 2.3.2 Matsuda’s Kummer-Artin-Schreier-Witt theory
$\operatorname{H}^{1}(\kappa,\mathbb{Z}/p^{s})\rightarrow\operatorname{H}^{1}(\mathcal{R}^{\rm bd},\mathbb{Z}/p^{s}).$ (15)
Theorem 2.7 ([4, Theorem 3.8]).
@@
$W_{s}(\mathcal{R}^{\operatorname{int}})$ ${\Psi_{s}}(\mathcal{R}^{\operatorname{int}})^{\times},$ (16)
$\underline{a}$ $\mapsto E_{s,p}(p^{n}\underline{a}).$
The morphism (16) effectively establishes a unification between Kummer and Artin-Schreier-Witt theories by relating a Witt vector $\underline{a}$ with a Kummer class $E_{s,p}(p^{n}\underline{a})$. Our strategy in this paper is to define the cover using Kummer theory while using an associated Witt vector to prove that the cover satisfies the desired conditions.
@@ 3.1 The Witt Vector Associated to a Kummer Extension
Given a $\mathbb{Z}/p^{s}$-cover of $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ described by the Kummer equation
$Z_{s}^{p^{s}}=\prod_{i=0}^{s-1}(u_{i})^{p^{i}}=:u\in\mathbb{K},$ (17)
where $\mathbb{K}$ is the function field of $\mathbb{P}^{s}_{R}$ and $u_{i}\in(\mathcal{R}^{\operatorname{int}})^{\times}\subset\mathbb{K}$. Let $\underline{X}^{\prime}=(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\in W_{s}(\mathcal{R}^{\operatorname{int}})$ be a Witt vector such that $\Psi_{s}(\underline{X}^{\prime})=u$, where $\Psi_{s}$ is Matsuda’s morphism (16). In this section, we prove the existence of the $u_{i}$ such that $\underline{X}^{\prime}\equiv(X_{0},\ldots,X_{s-1})\pmod{\pi}$.
@@
$E_{s,p}((0,\ldots,0,X_{0}^{\prime},\ldots,X_{m-1}^{\prime}))=E_{s-m,p}((X_{0}^{\prime},\ldots,X_{m-1}^{\prime})).$
@@
$X_{0}^{\prime}=\frac{1}{p\lambda}\ln(u_{0}),\;X_{1}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{u_{1}}{G_{2,p}((X_{0}^{\prime},0))}\right),\;\ldots,\;X_{s-1}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{u_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right).$
@@
Let us proceed by induction on $i$. Suppose first that $i=1$. Then it is necessary that $\Psi_{1}(X_{0}^{\prime})=\exp(p\lambda X_{0}^{\prime})=u_{0}$. Therefore, $X_{0}^{\prime}=\frac{1}{p\lambda}\ln(u_{0})$ as expected.
Assume the induction hypothesis for $s-1$. That is, one can find $X_{0}^{\prime},\ldots,X_{s-2}^{\prime}$ such that $E_{s-1,p}(p^{s-1}(X_{0}^{\prime},\ldots,X_{s-2}^{\prime}))=\prod_{i=0}^{s-2}(u_{i})^{p^{i}}$. We want to find $X_{s-1}^{\prime}$ such that
@@
Since we only need to find a solution, we take the $p^{s-1}$-th root of both sides, which gives
@@
The second equality follows from the decomposition of Witt vectors
$\underline{x}=(x_{0},\ldots,x_{s-1},x_{s},x_{s+1},\ldots)=(x_{0},\ldots,x_{s-1},0,0,\ldots)+(\underbrace{0,\ldots,0}_{s\text{ times}},x_{s},x_{s+1},\ldots)$
and the commutativity of $E_{s,p}$ by (13). Due to Proposition 3.1, we have
@@
3.2 Construct the Kummer Extension
Let us now further assume that
@@ 3.2 Construct the Kummer Extension
where $X_{0}^{\prime},\ldots,X_{i-1}^{\prime}$ are derived from $u_{0},\ldots u_{i-1}$ and Proposition 3.2.
@@
In the above settings, suppose $\nu(G_{i+1}^{\prime})>\frac{p}{p-1}$ for $i=0,\ldots,s$. Then $X_{i}^{\prime}\equiv X_{i}\mod(\zeta_{p^{i+1}}-1)$ for all $i$.
@@
The case $i=0$ is straightforward. Suppose the proposition holds for $i<s$. Applying Proposition 3.2, we obtain
$X_{s-1}^{\prime}$ $=\frac{1}{p\lambda}\ln\left(\frac{G_{s}(X_{0},\ldots,X_{s-2})+p\lambda X_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right)$
$=\frac{1}{p\lambda}\ln\left(1+\frac{-G^{\prime}_{s}+p\lambda X_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right)$
$=\frac{1}{p\lambda}\ln\left(1-G^{\prime}_{s}+p\lambda X_{s-1}+o\left(p^{p/(p-1)}\right)\right)$
$=X_{s-1}+\frac{1}{p\lambda}o\left(p^{p/(p-1)}\right),$
where $o\left(p^{p/(p-1)}\right)$ refers to terms with valuation strictly larger than $\frac{p}{p-1}$. The third equality follows from the fact that $\nu(G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))-1)\geq 1/(p^{s}(p-1))$ as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0}^{\prime},\allowbreak\ldots,X_{s-2}^{\prime}\rrbracket$ due to Lemma 2.4 and $X_{j}\equiv X^{\prime}_{j}\pmod{(\zeta_{p^{j+1}}-1)}$ for $j<s-1$ by induction, which allows us to write
$\frac{1}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}=1+O\left(p^{1/(p^{s}(p-1))}\right)$
as a series in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket X_{0},\allowbreak\ldots,X_{s-2}\rrbracket$. Here $O\left(p^{1/(p^{s}(p-1))}\right)$ refers to terms with valuation at least $1/(p^{s}(p-1))$. In addition, since $\nu(-G^{\prime}_{s}+p\lambda X_{s-1}+o\left(p^{p/(p-1)}\right))\geq\frac{p}{p-1}$, we have, for $m>1$,
$\nu\left(\frac{\left(-G^{\prime}_{s}+p\lambda X_{s-1}+o\left(p^{p/(p-1)}\right)\right)^{m}}{m}\right)>\frac{p}{p-1}.$
The final equality then follows from the series expansion of the $\ln(1+X)$ function. ∎
@@
For each $i$, one can inductively compute $G_{i+1}(X_{0},\ldots,X_{i-1})\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}]\allowbreak[X_{0},\ldots,X_{i-1}]$ such that:

### v3

@@ 2.3.1 Deformed Artin–Hasse exponential function
The Artin–Hasse exponential series at $p$ is the formal power series
@@
$E_{s,p}(t):=\frac{E_{p}(\zeta_{p^{s}}t)}{E_{p}(t)}=\exp\left(\sum_{i=0}^{s-1}(\zeta_{p^{s-i}}-1)\frac{t^{p^{i}}}{p^{i}}\right)\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket t\rrbracket,$ (12)
and for $\underline{a}=(a_{0},\ldots,a_{s-1})$,
$E_{s,p}(\underline{a}):=\prod_{i=0}^{s-1}E_{s-i,p}(a_{i})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket a_{0},\ldots,a_{s-1}\rrbracket,$ (13)
which can be rewritten as
$E_{s,p}(\underline{a})=\exp\left(\sum_{i=0}^{s-1}(\zeta_{p^{s-i}}-1)\frac{w_{i}}{p^{i}}\right),$ (14)
where $w_{i}=\sum_{j=0}^{i}p^{j}a_{j}^{p^{i-j}}$ is the $i$th ghost component of $\underline{a}$.
It follows from (14) that
@@
Define
$G_{s,p}(\underline{a}):=\frac{E_{s,p}(p\underline{a})}{E_{s-1,p}(\underline{a})}\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket a_{0},\dots,a_{s-1}\rrbracket.$
When $s=1$, we interpret $E_{0,p}=1$, so $G_{1,p}((a_{0}))=E_{1,p}(pa_{0})$.
By (15), we have
@@
We have
$G_{s,p}(\underline{a})\in\mathbb{Z}_{(p)}[\zeta_{p^{s}}]\llbracket(\zeta_{p^{s}}-1)a_{0},\dots,(\zeta_{p}-1)a_{s-1}\rrbracket.$
Proof.
Using the ghost-component form of $E_{m,p}$ and the definition $G_{s,p}(\underline{a})=E_{s,p}(p\underline{a})E_{s-1,p}(\underline{a})^{-1}$, we obtain
$G_{s,p}(\underline{a})$ $=\prod_{j=0}^{s-1}\exp\left(\sum_{i=0}^{s-j-1}\Big(p(\zeta_{p^{s-j-i}}-1)-(\zeta_{p^{s-j-1-i}}-1)\Big)\frac{a_{j}^{p^{i}}}{p^{i}}\right)$
$=\prod_{j=0}^{s-1}\exp\left(\sum_{i=0}^{s-j-1}(\zeta_{p^{s-j-i}}-1)\Big(p-(1+\zeta_{p^{s-j-i}}+\cdots+\zeta_{p^{s-j-i}}^{p-1})\Big)\frac{a_{j}^{p^{i}}}{p^{i}}\right)$
$=\prod_{j=0}^{s-1}E_{s-j,p}\left(\Big(p-1-[\zeta_{p^{s-j}}]-[\zeta_{p^{s-j}}]^{2}-\cdots-[\zeta_{p^{s-j}}]^{p-1}\Big)[a_{j}]\right),$
where $\cdot\,$] denotes the Teichmüller lift (in the Witt vectors of the appropriate length). Reducing modulo the ideal $(\zeta_{p^{s-j}}-1)$ gives $\zeta_{p^{s-j}}]\equiv 1$, hence
$p-1-[\zeta_{p^{s-j}}]-\cdots-[\zeta_{p^{s-j}}]^{p-1}\equiv p-p=0.$
By the standard identity
$\ker\big(W_{s-j}(A)\to W_{s-j}(A/I)\big)=W_{s-j}(I),$
valid for any ring $A$ and ideal $I\subset A$, $\ker\big(W_{s-j}(A)\to W_{s-j}(A/I)\big)=W_{s-j}(I)$, applied with $A=\mathbb{Z}_{(p)}[\zeta_{p^{s-j}}]$ and $I=(\zeta_{p^{s-j}}-1)$, it follows that
$p-1-[\zeta_{p^{s-j}}]-\cdots-[\zeta_{p^{s-j}}]^{p-1}\in W_{s-j}\big((\zeta_{p^{s-j}}-1)\mathbb{Z}_{(p)}[\zeta_{p^{s-j}}]\big).$
Therefore the Witt vector $\Big(p-1-[\zeta_{p^{s-j}}]-\cdots-[\zeta_{p^{s-j}}]^{p-1}\Big)[a_{j}]$ lies in the same Witt ideal, and the corresponding Artin–Hasse factor $E_{s-j,p}(\cdot)$ is a power series in the single variable $(\zeta_{p^{s-j}}-1)a_{j}$ with coefficients in $\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$. Taking the product over $j=0,\dots,s-1$ gives the stated inclusion. ∎
@@
The polynomials $E_{1},\ldots,E_{s}$ (resp. $G_{1},\ldots,G_{s}$) in Theorem 1.1 are approximations of $E_{1,p},\ldots,E_{s,p}$ (resp. $G_{1,p},\ldots,G_{s,p}$).
2.3.2 Matsuda’s Kummer–Artin–Schreier–Witt theory
Let $K$ be a discrete valuation field containing $\zeta_{p^{s}}$, with residue field $k$. The Robba ring $\mathcal{R}$ over $K$ is defined as the ring of Laurent series that converge in some annulus $\{z\mid\rho<\lvert z\rvert<1\}$ for some $0<\rho<1$. Let $\mathcal{R}^{\rm bd}$ be the subring of $\mathcal{R}$ consisting of formal sums with bounded coefficients. The ring $\mathcal{R}^{\rm bd}$ carries a Gauss norm:
@@ 2.3.2 Matsuda’s Kummer–Artin–Schreier–Witt theory
$\operatorname{H}^{1}(\kappa,\mathbb{Z}/p^{s})\longrightarrow\operatorname{H}^{1}(\mathcal{R}^{\rm bd},\mathbb{Z}/p^{s}).$ (17)
Theorem 2.8 ([14, Theorem 3.8]).
@@
[…]
$\underline{a}$ $\mapsto E_{s,p}(p^{s}\underline{a}).$
The homomorphism (18) establishes a link between Kummer and Artin–Schreier–Witt theories by associating to a Witt vector $\underline{a}$ the Kummer class $E_{s,p}(p^{s}\underline{a})$. Our strategy in this paper is to define the cover using Kummer theory while using an associated Witt vector to prove that the cover satisfies the desired conditions.
@@ 3.1 The Witt Vector Associated to a Kummer Extension
Recall that any $\mathbb{Z}/p^{s}$-cover of $\mathbb{P}^{s}_{R}=\operatorname{Proj}R[X_{0},\ldots,X_{s-1},Y]$ can be described by a Kummer equation of the form
$Z_{s-1}^{p^{s}}=\prod_{i=0}^{s-1}(u_{i})^{p^{i}}=:u\in K(X_{0},\ldots,X_{s-1}),$ (19)
where $K$ is the fraction field of $R$ and $k$ is the residue field of $K$.
Proposition 3.1.
There exists a Henselian local field $\mathbb{K}$ with residue field $k((x_{0},\ldots,x_{s-1}))$.
Proof.
We proceed by induction on $s\geq 1$. Here $k((x_{0},\ldots,x_{s-1}))$ denotes the iterated Laurent series field $k((x_{0}))\cdots((x_{s-1}))$.
For $s=1$, Lemma 2.7 applied with base field $K$ and Robba variable $X_{0}$ shows that $\mathcal{R}^{\mathrm{int}}_{K}(X_{0})$ is a Henselian DVR with residue field $k((x_{0}))$. Its fraction field
$\mathbb{K}_{1}:=\mathcal{R}^{\mathrm{bd}}_{K}(X_{0})$
is therefore a Henselian local field with residue $k((x_{0}))$.
Assume the statement holds for s1s-1, so there exists a Henselian local field $\mathbb{K}_{s-1}$ with residue field $k((x_{0},\ldots,x_{s-2}))$. Applying Lemma 2.7 again, now with base field $\mathbb{K}_{s-1}$ and a new Robba variable $X_{s-1}$, we find that $\mathcal{R}^{\mathrm{int}}_{\mathbb{K}_{s-1}}(X_{s-1})$ is a Henselian DVR whose residue field is
$\mathrm{res}\!\left(\mathcal{R}^{\mathrm{int}}_{\mathbb{K}_{s-1}}(X_{s-1})\right)=\mathrm{res}(\mathbb{K}_{s-1})((x_{s-1}))=k((x_{0},\ldots,x_{s-2}))((x_{s-1}))=k((x_{0},\ldots,x_{s-1})).$
Its fraction field
$\mathbb{K}_{s}:=\mathcal{R}^{\mathrm{bd}}_{\mathbb{K}_{s-1}}(X_{s-1})$
is then a Henselian local field with this residue.
By induction, the claim follows for all $s\geq 1$, with $\mathbb{K}:=\mathbb{K}_{s}$. ∎
For the remainder of this section, we set $\mathcal{R}^{\mathrm{bd}}=\mathbb{K}$ as in Proposition 3.1. Its associated Henselian DVR is $\mathcal{R}^{\mathrm{int}}=\mathcal{R}^{\mathrm{int}}_{\mathbb{K}_{s-1}}(X_{s-1})$, as in the proof of the same proposition, and its residue field is $k((x_{0},\ldots,x_{s-1}))$.
Lemma 3.2.
With the notation of Proposition 3.1, there is a canonical injective ring homomorphism
$R\llbracket X_{0},\ldots,X_{s-1}\rrbracket\hookrightarrow\mathcal{R}^{\mathrm{int}}_{\mathbb{K}_{s-1}}(X_{s-1})=\mathcal{R}^{\mathrm{int}}.$
In particular, any series in $R\llbracket X_{0},\ldots,X_{s-1}\rrbracket$ may be viewed as an element of the Henselian DVR $\mathcal{R}^{\mathrm{int}}$.
Proof.
We argue by induction on $s$.
If $s=1$, then $R\llbracket X_{0}\rrbracket\subset\mathcal{R}^{\mathrm{int}}_{K}(X_{0})$ because any series $f(X_{0})=\sum_{n\geq 0}a_{n}X_{0}^{n}$ with $a_{n}\in R$ has Gauss norm $\max_{n}|a_{n}|\leq 1$, hence lies in the integral Robba ring. Injectivity follows because the Robba variable $X_{0}$ is transcendental over $K$.
Assume $s>1$ and that we already have a canonical embedding $R\llbracket X_{0},\ldots,X_{s-2}\rrbracket\hookrightarrow\mathcal{R}^{\mathrm{int}}_{\mathbb{K}_{s-2}}(X_{s-2})=\mathcal{O}_{\mathbb{K}_{s-1}}$. Now take $f\in R\llbracket X_{0},\ldots,X_{s-1}\rrbracket$ and write it uniquely as a power series in $X_{s-1}$,
$f=\sum_{n\geq 0}a_{n}\,X_{s-1}^{n},\qquad a_{n}\in R\llbracket X_{0},\ldots,X_{s-2}\rrbracket.$
Via the inductive embedding, each coefficient $a_{n}$ defines an element of $\mathcal{O}_{\mathbb{K}_{s-1}}$. Hence the same formula defines an element of $\mathcal{O}_{\mathbb{K}_{s-1}}\llbracket X_{s-1}\rrbracket$, which is contained in $\mathcal{R}^{\mathrm{int}}_{\mathbb{K}_{s-1}}(X_{s-1})$ by the same Gauss-norm argument as for $s=1$. This gives the desired map; injectivity again follows from the transcendence of $X_{s-1}$ over $\mathbb{K}_{s-1}$. ∎
To construct the isogeny $\mathcal{W}_{s}\rightarrow\mathcal{V}_{s}$ in (2), which is generically an endomorphism of $\mathbb{G}_{m}$, we consider a $\mathbb{Z}/p^{s}$-cover of $\mathbb{P}^{s}_{R}$ defined by a Kummer equation of the form (19). We require that this cover has good reduction and that its special fiber is a $\mathbb{Z}/p^{s}$-cover of $\mathbb{P}^{s}_{k}$ associated with the Witt vector $(x_{0},\ldots,x_{s-1})$. Assume that the cover has separable reduction, and let $\underline{X}^{\prime}=(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})\in W_{s}(\mathcal{R}^{\mathrm{int}})$ be a Witt vector such that $\Psi_{s}(\underline{X}^{\prime})=u$, where $\Psi_{s}$ is Matsuda’s morphism (18). Our task is to prove the existence of such a $u$ for which, for each $i=0,1,\ldots,s-1$, the component $X_{i}^{\prime}$ is congruent to $x_{i}$ modulo the maximal ideal of $R$.
@@
$E_{s,p}((0,\ldots,0,X_{0}^{\prime},\ldots,X_{m-1}^{\prime}))=E_{m,p}((X_{0}^{\prime},\ldots,X_{m-1}^{\prime})).$
@@
$X_{0}^{\prime}=\frac{1}{p\lambda}\ln(u_{0}),\quad X_{1}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{u_{1}}{G_{2,p}((X_{0}^{\prime},0))}\right),\ldots,\quad X_{s-1}^{\prime}=\frac{1}{p\lambda}\ln\left(\frac{u_{s-1}}{G_{s,p}((X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0))}\right),$
provided the $\ln$ function is well-defined.
@@
We proceed by induction on $s$. Suppose first that $s=1$. Then it is necessary that $\Psi_{1}(X_{0}^{\prime})=\exp(p\lambda X_{0}^{\prime})=u_{0}$. Therefore, $X_{0}^{\prime}=\frac{1}{p\lambda}\ln(u_{0})$, as expected.
Assume the induction hypothesis for s1s-1. That is, one can find $X_{0}^{\prime},\ldots,X_{s-2}^{\prime}$ such that
$E_{s-1,p}(p^{s-1}(X_{0}^{\prime},\ldots,X_{s-2}^{\prime}))=\prod_{i=0}^{s-2}(u_{i})^{p^{i}}.$
We want to find $X_{s-1}^{\prime}$ such that
@@
By (14), for any Witt vector $\underline{a}$ of length $m$ and any $N\geq 0$, we have
$E_{m,p}(p^{N}\underline{a})=E_{m,p}(\underline{a})^{p^{N}}.$
Hence
$E_{s,p}(p^{s}\underline{X}^{\prime})=E_{s,p}(p\underline{X}^{\prime})^{p^{s-1}}\quad\text{and}\quad E_{s-1,p}(p^{s-1}\underline{X}^{\prime}_{<s-1})=E_{s-1,p}(\underline{X}^{\prime}_{<s-1})^{p^{s-1}}.$
Therefore, to produce a solution at level $p^{s}$, it suffices to solve
@@
The second equality follows from the decomposition
$(X_{0}^{\prime},\ldots,X_{s-1}^{\prime})=(X_{0}^{\prime},\ldots,X_{s-2}^{\prime},0)+(0,\ldots,0,X_{s-1}^{\prime})$
and the commutativity of $E_{s,p}$ in (15). By Proposition 3.3, we have
@@
3.2 Constructing the Kummer Extension
Assume further that
@@ 3.2 Constructing the Kummer Extension
where $X_{0}^{\prime},\ldots,X_{i-1}^{\prime}$ are derived from $u_{0},\ldots,u_{i-1}$ and Proposition 3.4.
@@
In the above setting, suppose $\nu(G_{i+1}^{\prime})>\frac{p}{p-1}$ for $i=0,\ldots,s-1$. Then $X_{i}^{\prime}\equiv X_{i}\mod(\zeta_{p^{s}}-1)$ for all $i$.
@@
We argue by induction on $i=0,\ldots,s-1$. We write $O(p^{a})$ (resp. $o(p^{a})$) for an element of valuation $\geq a$ (resp. >a>a).
For $i=0$, we have $u_{0}=G_{1}+p\lambda X_{0}=1+p\lambda X_{0}$ and hence
$X_{0}^{\prime}=\frac{1}{p\lambda}\ln(u_{0})=\frac{1}{p\lambda}\ln(1+p\lambda X_{0})=X_{0}+\frac{1}{p\lambda}\,o\bigl(p^{p/(p-1)}\bigr),$
so $\nu(X_{0}^{\prime}-X_{0})>0$. Since $R=\mathbb{Z}_{(p)}[\zeta_{p^{s}}]$ is a DVR and $\nu(\zeta_{p^{s}}-1)=\frac{1}{p^{s-1}(p-1)}$ is its smallest positive valuation, this implies $X_{0}^{\prime}\equiv X_{0}\pmod{(\zeta_{p^{s}}-1)}$.
Assume now $1\leq i\leq s-1$ and that $X_{j}^{\prime}\equiv X_{j}\pmod{(\zeta_{p^{s}}-1)}$ for all $j<i$. By Proposition 3.4 and the definition of $u_{i}$ in (20), we have
$X_{i}^{\prime}$ $=\frac{1}{p\lambda}\ln\left(\frac{u_{i}}{G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))}\right)=\frac{1}{p\lambda}\ln\left(\frac{G_{i+1}(X_{0},\ldots,X_{i-1})+p\lambda X_{i}}{G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))}\right)$
$=\frac{1}{p\lambda}\ln\left(1+\frac{-G_{i+1}^{\prime}+p\lambda X_{i}}{G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))}\right).$
Set $D_{i}:=G_{i+1,p}((X_{0}^{\prime},\ldots,X_{i-1}^{\prime},0))$. By Lemma 2.5, we have $\nu(D_{i}-1)\geq\nu(\zeta_{p^{i+1}}-1)$, so $D_{i}^{-1}=1+O\bigl(p^{\nu(\zeta_{p^{i+1}}-1)}\bigr)$. Using the hypothesis $\nu(G_{i+1}^{\prime})>\frac{p}{p-1}$, we get
$\frac{-G_{i+1}^{\prime}+p\lambda X_{i}}{D_{i}}=-G_{i+1}^{\prime}+p\lambda X_{i}+o\bigl(p^{p/(p-1)}\bigr).$
Consequently,
$X_{i}^{\prime}=\frac{1}{p\lambda}\ln\left(1+p\lambda X_{i}-G_{i+1}^{\prime}+o\bigl(p^{p/(p-1)}\bigr)\right)=X_{i}+\frac{1}{p\lambda}\,o\bigl(p^{p/(p-1)}\bigr),$
where the last equality uses the power-series expansion of $\ln(1+T)$ and the fact that for $m>1$ one has $\nu(T^{m}/m)>\frac{p}{p-1}$ as soon as $\nu(T)\geq\frac{p}{p-1}$. Thus $\nu(X_{i}^{\prime}-X_{i})>0$, so by discreteness of the valuation on $R$ we conclude $X_{i}^{\prime}\equiv X_{i}\pmod{(\zeta_{p^{s}}-1)}$. ∎
@@
The following lemma proves the stability of the Tate condition under substitution.
Lemma 3.7.
Let $R$ be a discretely valued ring with valuation $\nu$. Let $f\in R\llbracket T_{1},\ldots,T_{m}\rrbracket$ satisfy the Tate condition. Suppose $g_{1},\ldots,g_{m}\in R\llbracket X_{1},\ldots,X_{n}\rrbracket$ satisfy the Tate condition and have zero constant terms. Then $f(g_{1},\ldots,g_{m})\in R\llbracket X_{1},\ldots,X_{n}\rrbracket$ satisfies the Tate condition.
Proof.
Write
$f=\sum_{\alpha\in\mathbb{Z}_{\geq 0}^{m}}a_{\alpha}T^{\alpha},\qquad T^{\alpha}=T_{1}^{\alpha_{1}}\cdots T_{m}^{\alpha_{m}},\qquad|\alpha|:=\alpha_{1}+\cdots+\alpha_{m}.$
Since each $g_{j}$ has zero constant term, we have $g_{j}\in(X_{1},\ldots,X_{n})$, hence
$(g_{1},\ldots,g_{m})^{\alpha}=g_{1}^{\alpha_{1}}\cdots g_{m}^{\alpha_{m}}\in(X_{1},\ldots,X_{n})^{|\alpha|}.$
Fix a multi-index $\beta$ and write $c_{\beta}\in R$ for the coefficient of $X^{\beta}$ in $f(g_{1},\ldots,g_{m})$. Since $(g_{1},\ldots,g_{m})^{\alpha}$ has total degree $\geq|\alpha|$, only indices $\alpha$ with $|\alpha|\leq|\beta|$ can contribute to $c_{\beta}$.
Now fix $M\in\mathbb{R}$. By the Tate condition on $f$, there exists $D$ such that $\nu(a_{\alpha})>M$ for all $\alpha$ with $|\alpha|>D$. Decompose
$c_{\beta}=\sum_{|\alpha|\leq D}a_{\alpha}\cdot\bigl[X^{\beta}\bigr](g_{1},\ldots,g_{m})^{\alpha}\;+\sum_{D<|\alpha|\leq|\beta|}a_{\alpha}\cdot\bigl[X^{\beta}\bigr](g_{1},\ldots,g_{m})^{\alpha},$
where $X^{\beta}](\cdots)$ denotes “coefficient of $X^{\beta}$”.
(1) The contribution from $|\alpha|>D$. Every term in the second sum has valuation >M>M because it is multiplied by $a_{\alpha}$.
(2) The contribution from $|\alpha|\leq D$. There are only finitely many such $\alpha$. Fix one. When we expand $(g_{1},\ldots,g_{m})^{\alpha}=g_{1}^{\alpha_{1}}\cdots g_{m}^{\alpha_{m}}$, the coefficient of $X^{\beta}$ is a finite sum of products of $|\alpha|$ coefficients taken from the power series $g_{1},\ldots,g_{m}$, where the total degrees of the chosen monomials add up to $|\beta|$. Therefore at least one of these chosen monomials has total degree $\geq|\beta|/|\alpha|$.
Since each $g_{j}$ satisfies the Tate condition, the valuations of coefficients of $g_{j}$ tend to $\infty$ as the total degree tends to $\infty$. Hence, for fixed $\alpha$, we have
$\nu\!\left(\bigl[X^{\beta}\bigr](g_{1},\ldots,g_{m})^{\alpha}\right)\longrightarrow\infty\qquad\text{as }|\beta|\to\infty.$
Multiplying by the fixed scalar $a_{\alpha}$ does not affect the conclusion.
Combining (1) and (2), we see that for every $M$ there exists $B$ such that $\nu(c_{\beta})>M$ for all $|\beta|>B$. This is precisely the Tate condition for $f(g_{1},\ldots,g_{m})$. ∎
@@
For each $i$, one can compute $G_{i+1}(X_{0},\ldots,X_{i-1})\in\mathbb{Z}_{(p)}[\zeta_{p^{i+1}}][X_{0},\ldots,\allowbreak X_{i-1}]$ inductively such that: